OpenStudy (anonymous):

Solve for the unknown. 2/3x-1 + 3/3x+1 = 4

OpenStudy (lilg132):

2/3x + 3/3x +1 -1 = 4 2/3x + 3/3x = 4 5/3x = 4 5/3x / (5/3) = 4 / (5/3) x = 2.4

OpenStudy (lilg132):

dont forget to click good answer :D

OpenStudy (anonymous):

i think i wrote the wrong equation. it should be like this. 2/(3x-1) + 3/(3x+1) = 4 my instructor told me to use quadratic formula here but i dont know how. :D

OpenStudy (lilg132):

the formula to use is